OpenStudy (anonymous):

I will give u a medal if u answer the question correctly and show work Solve the problem using Cramers Rule. Ana has a piggy bank that contains nickels, dimes, and quarters. When she opens the bank and counts her money, she has $3.35. She counted twice as many nickels as quarters, and two more dimes than quarters. How many dimes does she have?

OpenStudy (anonymous):

Here is the answer with show work. 5n+10d+25q = 335 n=2q d=q+2 Standard form: 10d+5n+25q=335; n-2q=0; d-q=2; substitute/eliminate d = -1/2n-5/2q+67/2 substitute/eliminate n = +2q-0 substitute/eliminate q = +7 Solution: d = 9; n = 14; q = 7;